Fermi National Laboratory


Accelerator Update

Wednesday December 17
The day shift began with an access into the Tevatron (TeV) sector D4, with Operations monitoring the largest Antiproton Source (Pbar) stack ever achieved of 245.017 mA, and with beam to MiniBooNE.

At 8 AM, a phase problem with a Linac RF station (KRF1) held off Linac beam. While an expert investigated the phase came back to its normal range.

Technicians came out of the TeV access at 9 AM.

At 9:24 AM, high voltage techs racked feeder 23 back into service.

Operations put the TeV through a dry squeeze at 10:52 AM and then put the TeV into shot setup at 1 PM.

At 2:57 PM, Operations established store 3101 with an initial luminosity of 2.8E31. Pbar resumed stacking.

At 7:49 PM, Operations stopped the Main Injector (MI) ramp so Recycler experts could conduct studies. MI went back to normal ramping at 8:08 PM. Recycler then resumed the study at 9:10 PM and ran for about an hour. Stacking resumed at 8:17 PM.

Thursday December 18
The midnight shift began with Operations monitoring store 3101, with Pbar stacking, and with MiniBooNE taking beam.

At 5 AM, Recycler experts began a study.

At 9 AM, Recycler began sending shots of antiprotons to MI. This took about two hours. Stacking was on hold during this time.

Operations terminated store 3101 at 12:12 PM.

Operations put the TeV into shot setup at 12:50 PM.

At 1:27 PM, a problem caused nearly all the MI RF station power supplies to trip off. An expert had the problem resolved by 1:48 PM.

Operations established store 3103 at 3:10 PM with a RECORD luminosity of 49.92E30. Then MI and Switchyard experts began a Switchyard (SY) study period.

Stacking resumed after the MI/SY study ended at 3:53 PM.

A Booster expert began a study period at 4:18 PM. Stacking resumed at 4:48 PM.

At 8:50 PM, Recycler finished their studies.

Friday December 19
The day shift began with Operations monitoring store 3103, with Pbar stacking, and with MiniBooNE taking beam.

At 2:45 AM, the TeV sector E1 wet engine alarmed on a gas pressure alarm. Operators monitored the system. Later, the bottle that supplies this gas was replaced.

Booster studies ended at 3:16 AM.

A Linac expert conducted studies from 4:12 AM to 5:53 AM.

A Recycler expert began a study at 4:50 AM.

Weekend Plans
The plan for today and this weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/19/2003   email Fermilab